OpenStudy (anonymous):

Help!!! will give medal :) The perimeter of a t.v is 150 inches. The ratio of the length to the width is 16 to 9. Find the length and width

OpenStudy (anonymous):

Let the length=16x width=9 x P=2(l+w) calculate x

OpenStudy (anonymous):

3=x is what i get

OpenStudy (anonymous):

find length and with

OpenStudy (anonymous):

so then do i just take 3 times 16 and 3 times 9?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

???

OpenStudy (anonymous):

48,27

OpenStudy (anonymous):

Ok thanks for the help:)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Another equation formulation: \[\left\{\frac{16}{9}=\frac{L}{w},L+w=75\right\} \]\[\{L\to 48,w\to 27\} \]